# MEEG and AT-DGNN: Advancing EEG Emotion Recognition with Music and Graph Learning

### v1

@@ Abstract
Recent advances in neuroscience have elucidated the crucial role of coordinated brain region activities during cognitive tasks. To explore this complexity, we introduce the MEEG dataset—a comprehensive multi-modal music-induced electroencephalogram (EEG) dataset and the Attention-based Temporal learner with Dynamic Graph Neural Network (AT-DGNN), a novel framework for EEG-based emotion recognition. The MEEG dataset captures a wide range of emotional responses to music, enabling an in-depth analysis of brainwave patterns in musical contexts. The AT-DGNN combines an attention-based temporal learner with a dynamic graph neural network (DGNN) to accurately model the local and global graph dynamics of EEG data across varying brain network topologies. Our evaluations show that AT-DGNN achieves superior performance, with accuracies (ACC) of 83.06% in arousal and 85.31% in valence, outperforming state-of-the-art (SOTA) methods on the MEEG dataset. Comparative analyses with traditional datasets like DEAP highlight the effectiveness of our approach and underscore the potential of music as a powerful medium for emotion induction. This study not only advances our understanding of the brain’s emotional processing but also enhances the accurcy of emotion recognition technologies in brain-computer interfaces (BCI), leveraging both graph-based learning and the emotional impact of music. The source code and dataset are available at https://github.com/xmh1011/AT-DGNN.
Index Terms—Electroencephalogram (EEG), Emotion recognition, Dynamic graph neural networks (DGNNs).
@@ I Introduction
Emotion, a complex physiological and psychological phenomenon, plays a critical role in interpreting genuine reactions to daily interactions [1]. In 1980, Russell introduced the dimensional models of emotion, with the Valence-Arousal (VA) model being particularly noteworthy [2]. These models conceptualize emotions within a continuous dimensional space, providing a framework that allows for a more nuanced quantification of emotional states.
@@
BCI technology, which decodes brain signals into commands for external devices, predominantly employs EEG signals [3]. This technology processes EEG signals through stages of preprocessing, feature extraction, and classification, thereby translating thoughts into actionable commands and enhancing human-machine interaction [4]. EEG signals, reflecting cortical voltage fluctuations, are crucial for emotion recognition due to their objectivity, high temporal resolution, affordability, and rapid data acquisition [5].
@@
EEG signals have long been pivotal in neuroscience research [6]. In 2011, Koelstra et al. established the DEAP emotion dataset by collecting physiological signals associated with emotional responses [7]. Subsequently, in 2015, Zheng et al. developed the SEED dataset, providing well-annotated EEG signals from multiple subjects [8]. Both datasets have been instrumental in facilitating the development and evaluation of computational models for emotion analysis. Over recent decades, diverse machine learning and signal processing techniques have been utilized for emotion recognition from physiological signals.
@@
A significant advancement came in 2018 when Lawhern et al. introduced EEGNet, a compact convolutional neural network (CNN) designed for efficient EEG signal processing [9]. Further advancements were made by Schirrmeister et al., who enhanced EEG decoding using deep learning with CNNs and provided sophisticated visualization tools to better comprehend brain activities [10]. Temporal convolutional networks (TCNs) have also become a formidable alternative for real-time BCI applications. In 2020, Ingolfsson et al. developed EEG-TCNet, demonstrating high accuracy in motor imagery tasks and suggesting that TCNs could surpass traditional methods in specific BCI applications [11]. This was supported by Musallam et al., who highlighted the versatility and efficiency of TCNs in complex BCI tasks through the fusion of TCN for motor imagery classification [12].
@@
All the aforementioned research treated EEG signals as two-dimensional time-series data, with the dimensions being channels and time [13]. Channels refer to EEG electrodes, which are strategically placed on the scalp to record brain activity from various functional areas according to the 10-20 system established by the International Federation of Clinical Neurophysiology [3]. In recent years, EEG data has been treated as a graph where electrodes are positioned in a two-dimensional layout that reflects their physical placement on the scalp [14]. In this context, the raw data or extracted features from each electrode introduce a third dimension to this spatial map [15].
In 2023, Ding et al. [16] introduced the Local-Global-Graph network (LGGNet), which processed EEG data in an image-like format and has shown enhanced classification performance, although it was primarily designed for multi-task classification and didn’t significantly improve emotion recognition accuracy compared to other models. In terms of EEG datasets, the availability of open-source, multi-modal EEG emotion datasets remains limited, with DEAP and SEED being some of the few prominent resources available.
@@
To address limitations in emotion recognition research, this study introduces the MEEG dataset, a multi-modal EEG emotion dataset inspired by the DEAP format but enhanced with music-induced emotional states for greater validity. Additionally, the LGGNet framework has been refined by integrating an attention mechanism and utilizing DGNN, significantly improving classification accuracy on the MEEG dataset. This approach facilitates a deeper exploration of interconnections between various brain functions through dynamic graph analysis.
The contributions of this study are threefold and can be summarized as follows:
@@
    We establish the MEEG dataset, a robust multi-modal EEG emotion dataset similar to DEAP but enhanced with music stimuli to effectively induce emotional states. This enhancement ensures higher ACC and F1 scores compared to DEAP using the same models.
@@
    We propose the AT-DGNN framework for exploring connections within and between different functional areas of the brain. By integrating an attention mechanism and employing DGNN, the LGGNet architecture is significantly enhanced for emotion analysis tasks.
@@ II Related Work
Figure 1: Structure of AT-DGNN. AT-DGNN model consists of two core modules: a feature extraction module and a dynamic graph neural network learning module. The feature extraction module (a) includes a temporal learner, multi-head attention mechanism, and a temporal convolution module. These components effectively harness local features of EEG signals through a sliding window technique, thereby enhancing the model’s ability to dynamically extract complex temporal patterns in EEG signals. In the graph learning module (b), the model first employs local filtering layers to segment and filter features from specific brain regions. Subsequently, it utilizes stacked dynamic graph convolutions to learn the intricate interactions between different brain regions, further enhancing AT-DGNN’s capability to integrate spatiotemporal features.
@@ II-A Multi-head Attention Mechanism
Each self-attention head comprises three main components: queries ($Q$), keys ($K$), and values ($V$). These components interact to generate attention scores that determine the weighted values.The procedure begins with the normalization of the input $X_{w}$ through a LayerNorm layer, a critical step for stabilizing the learning process:
$q_{ht}=k_{ht}=v_{ht}=W^{Q}\text{LN}(X_{w,t})$ (1)
This formulation demonstrates how queries, keys, and values are generated from the normalized data at each position $t$ for each head $h$, thus aiding in the computation of attention scores.
The context vector for each head is computed as a weighted sum of the values, where weights are determined by softmax-normalized alignment scores. These scores are derived using the scaled dot-product attention mechanism:
@@
Ultimately, the outputs from all the heads are concatenated and subjected to a linear transformation. This consolidated output of the MHA layer is then merged with the initial input via a residual connection and normalized, leading to an optimized attention output from the layer.
@@ II-B Graph Neural Networks
Graph neural networks (GNNs) represent a significant advancement in the domain of neural networks, designed explicitly to process graph-structured data. Unlike CNNs, GNNs excel in capturing the intricate relationships and dependencies among nodes within a graph through processes of aggregation and propagation of information across local neighbors [17]. A typical graph is denoted as $G=(V,E)$, where $V$ symbolizes the set of nodes and $E$ the set of edges. Each node $v_{i}\in V$ and edge $e_{ij}=(v_{i},v_{j})\in E$ can be respectively associated with a node and an edge in the graph. The adjacency matrix $A$ is configured as an $n\times n$ matrix where $A_{ij}=1$ if $e_{ij}\in E$ and $A_{ij}=0$ otherwise. Node attributes are represented by $X$, where $X\in\mathbb{R}^{n\times d}$, with each $x_{i}\in\mathbb{R}^{d}$ denoting the feature vector of node $v_{i}$.
@@
$h_{t,i}^{(k+1)}=\sigma\left(\sum_{j\in\mathcal{N}(v_{i})}\alpha_{t,ij}^{(k)}W^{(k)}h_{t,j}^{(k)}+b^{(k)}\right)$ (3)
Here, $\mathcal{N}(v_{i})$ represents the neighbors of node $v_{i}$, $\alpha_{t,ij}^{(k)}$ are the attention coefficients indicating the significance of the features of neighbor $j$ to node $i$, and $W^{(k)}$ and $b^{(k)}$ are trainable parameters of the $k$-th layer, with $\sigma$ being a non-linear activation function.
@@ II-C Graph Neural Networks for EEG
In 2019, Song et al. [18] first introduced a novel DGCNN specifically for multichannel EEG emotion recognition, employing an adjacency matrix to dynamically model EEG channel relationships and enhance feature discrimination. This method proved to be superior to existing approaches. Subsequently, advancements in GNNs have been increasingly applied to emotion recognition. On the SEED dataset, Bao et al. [19] integrated a multi-layer GNN with a style-reconfigurable CNN, while Asadzadeh et al. [20] further improved the DGCNN by incorporating Bayesian signal recovery techniques, both achieving enhanced performance. For the DEAP dataset, Zhao et al. [21] developed coherence-based node clustering and multi-pooling techniques for GCNs. Additionally, Zhang et al. [22] enhanced GCN feature extraction by using Granger causality and differential entropy to construct more informative connectivity matrices.
@@ III Methodology
In this section, we will introduce the detailed architecture of AT-DGNN from the three aspects of data preprocessing, feature extraction and graph learning in turn, and the overall structure of AT-DGNN is shown in Fig 1.
@@ III-A EEG Data Preprocessing
The initial sampling rate of MEEG dataset is 1000 Hz. However, since most existing EEG datasets are sampled at either 128 Hz or 200 Hz, the data are downsampled to 200 Hz. This downsampling is crucial because the original EEG data often present significant challenges, including high levels of artifact noise, low signal amplitudes, poor interference resistance, a limited frequency range, and overlapping interference signals within the EEG band. Consequently, preprocessing of the raw EEG data is necessary before feature extraction can commence. Typically, EEG signals are categorized into five distinct frequency bands: Delta (1-4 Hz), Theta (4-8 Hz), Alpha (8-14 Hz), Beta (14-31 Hz), and Gamma (31-50 Hz) [23]. Accordingly, a band-pass filter ranging from 1 to 50 Hz is applied to the EEG data to eliminate interference from extraneous factors.
@@ III-B1 Temporal Learner
We introduce a temporal learner layer employing multiscale 1D temporal kernels ($T$ kernels) that directly extract dynamic temporal representations from EEG data $X_{i}\in\mathbb{R}^{E\times T}$, where $E$ denotes the number of EEG electrodes, and $T$ is the sample length. These kernels obviate the need for manually extracted features. The kernels’ length, dictated by the EEG data’s sampling frequency ($f_{s}$) and scaling coefficients ($\alpha^{i}$), is given by:
@@
EEG data processed through these layers yield dynamic time-frequency representations. An average pooling (AvgPool) layer, acting as a window function, calculates the averaged power across shorter segments. Applying logarithmic transformations and squaring functions further refines the feature extraction, as described by Schirrmeister et al. [10]. The output from each layer $i$, denoted as $Z^{i}_{\text{temp}}$, is formulated as:
$Z^{i}_{\text{temp}}=\Phi_{\text{log}}(\text{AvgPool}(\Phi_{\text{square}}(\mathcal{F}_{\text{Conv1-D}}(X_{i},S^{i}_{T}))))$ (5)
Here, $S^{i}_{T}$ indicates the kernel size, $\Phi_{\text{log}}$ is the logarithmic activation function, $\Phi_{\text{square}}$ represents the square function, and $\mathcal{F}_{\text{Conv1-D}}$ signifies the 1D convolution operation.
@@
$Z_{T}=f_{\text{bn}}(Z^{1}_{\text{temp}},Z^{2}_{\text{temp}},Z^{3}_{\text{temp}})$ (6)
where $f_{\text{bn}}$ denotes the batch normalization operation.
@@ III-B2 Sliding Window Segmentation
After temporal learner, we utilize a sliding window approach to segment the time series into multiple windows, which effectively extracts local features of EEG signals and augments the data. To mitigate the computational overhead introduced by the sliding window step, we employ the convolution-based sliding window method proposed by Schirrmeister et al[24]. This method integrates the sliding window approach with convolutional layers, allowing convolution operations to be performed only once across all windows, thereby reducing computation and inference time through parallel processing.
We employ a sliding window with a length of $W$ and a stride of $S$ to segment the time series $X$ into multiple windows $X_{w}\in\mathbb{R}^{B\times C\times W}$, where $w=1,\ldots,n$ denotes the window index and $n$ is the total number of windows. Each window $X_{w}$ is processed sequentially through subsequent attention and temporal convolution blocks. The formula for calculating the number of windows $n$ is:
$n=\left\lfloor\frac{\text{length}-\text{window\_size}}{\text{stride}}\right\rfloor+1$ (7)
where length is the length of the input sequence, window_size is the length of each window, and stride is the step size.

### v2

# MEEG and AT-DGNN: Improving EEG Emotion Recognition with Music Introducing and Graph-based Learning

@@ Abstract
Neuropsychological research highlights the essential role of coordinated activities across brain regions during cognitive tasks. This study expands the existing EEG datasets by constructing the MEEG dataset, a multi-modal compilation of music-induced electroencephalogram (EEG) recordings, and further investigates the brain network topology during emotional responses to music. The MEEG dataset, capturing emotional responses to various musical stimuli across different valence and arousal levels, enables an in-depth analysis of brainwave patterns within musical contexts. We introduce the Attention-based Temporal Learner with Dynamic Graph Neural Network (AT-DGNN), a novel framework for EEG emotion recognition. By integrating an attention mechanism with a dynamic graph neural network (DGNN), the AT-DGNN model captures complex local and global EEG dynamics, demonstrating superior performance with accuracy of 83.74% in arousal and 86.01% in valence, outperforming current state-of-the-art (SOTA) methods. Comparative analyses with traditional datasets like DEAP underscore our model’s effectiveness, highlighting the potential of music as a potent emotional stimulus. This study advances graph-based learning techniques in brain-computer interfaces (BCI), significantly enhancing the precision of EEG-based emotion recognition and deepening our understanding of cognitive functions in various brain regions. The source code and dataset are accessible at https://github.com/xmh1011/AT-DGNN.
Index Terms—electroencephalogram, emotion recognition, dynamic graph neural networks.
@@ I Introduction
Emotion, a complex physiological and psychological phenomenon, plays a critical role in interpreting genuine reactions to daily interactions [1]. In 1980, Russell introduced the dimensional models of emotion, with the Valence-Arousal (VA) model being particularly noteworthy [2]. The model conceptualize emotions within a continuous dimensional space, providing a framework that allows for a more nuanced quantification of emotional states.
@@
BCI technology decodes EEG signals, which reflect cortical voltage fluctuations, into commands for external devices, enhancing human-machine interaction [3, 4]. This technology involves preprocessing, feature extraction, and classification stages, translating thoughts into actionable commands. Due to their objectivity, high temporal resolution, affordability, and rapid acquisition, EEG signals are indispensable for emotion recognition and neuroscience research [5, 6].
@@
In 2011, Koelstra et al. established the DEAP dataset, which consists of physiological signals, such as EEG, elicited by emotional responses to musical videos [7]. Subsequently, Zheng et al. introduced the SEED dataset in 2015, comprising well-annotated EEG signals induced by cinematic stimuli, with data collected from multiple subjects [8]. These datasets have significantly contributed to the advancement of computational models for emotion analysis. Over the past decades, a variety of machine learning and signal processing techniques have been applied to emotion recognition based on physiological data.
@@
Recent advances in deep learning have significantly impacted EEG signal processing. In 2018, Lawhern et al. proposed EEGNet, a compact convolutional neural network (CNN) tailored for EEG analysis [9]. Subsequent developments by Schirrmeister et al. further refined EEG decoding capabilities using deep CNN architectures, introducing sophisticated visualization tools [10]. The adoption of temporal convolutional networks (TCNs) has introduced robust alternatives for real-time BCI applications. Notably, in 2020, Ingolfsson et al. developed EEG-TCNet, which demonstrated superior accuracy in motor imagery tasks, suggesting that TCNs could outperform traditional methods in specific BCI contexts [11]. This assertion was corroborated by Musallam et al., who demonstrated the versatility and efficiency of TCNs in complex BCI tasks, particularly through the integration of TCNs in motor imagery classification [12].
@@
All the aforementioned research treated EEG signals primarily as two-dimensional time-series data, with dimensions corresponding to channels and time [13]. Here, channels denote EEG electrodes, positioned on the scalp using the 10-20 system, as prescribed by the International Federation of Clinical Neurophysiology, to capture neural activity from distinct brain regions [3]. More recently, a paradigm shift has occurred with EEG data now being conceptualized as graphs, representing the electrodes’ two-dimensional arrangement on the scalp [14]. In this model, raw data or extracted features from each electrode augment this spatial framework by adding a third dimension [15].
In 2023, Ding et al. introduced the Local-Global-Graph network (LGGNet) to process EEG data in an image-like format, improving classification performance [16]. However, LGGNet’s single-layer graph neural network (GNN) struggles to capture the dynamic features of EEG signals. Additionally, popular EEG datasets like DEAP and SEED primarily use video stimuli, limiting the study of the relationship between music, emotions, and EEG signals. In contrast, EEG datasets and research using other types of stimuli are relatively scarce.
@@
To address shortcomings in emotion recognition research, we create the MEEG dataset, a multi-modal EEG emotion dataset in the DEAP format with music-induced emotional states to enhance ecological validity. We improve the LGGNet architecture by integrating an attention mechanism and using a multi-layer dynamic graph neural network, significantly boosting classification accuracy on the MEEG dataset. This innovative approach allows for a more nuanced examination of brain function interconnections via graph-based learning.
The contributions of our study are threefold and can be summarized as follows:
@@
    We establish the MEEG dataset, a multi-modal EEG emotion dataset in the DEAP format, enhanced with various types of music to effectively induce emotional states. Experimental results show that the MEEG dataset outperforms the video-stimuli-based DEAP dataset in emotion induction, effectively enhancing model ACC and F1 scores.
@@
    We propose the AT-DGNN framework for exploring connections within and between different functional areas of the brain. By integrating an attention mechanism and employing DGNN, the LGGNet architecture is significantly enhanced for emotion recognition.
@@ II Related Work
Figure 1: Structure of AT-DGNN. The AT-DGNN model comprises two core modules: a feature extraction module (a) and a dynamic graph neural network learning module (b). The feature extraction module consists of a temporal learner, a multi-head attention mechanism, and a temporal convolution module. These components effectively leverage local features of EEG signals through a sliding window technique, thereby enhancing the model’s capacity to dynamically extract complex temporal patterns in EEG signals. In the graph-based learning module, the model initially employs local filtering layers to segment and filter features from specific brain regions. Subsequently, the architecture employs three layers of stacked dynamic graph convolutions to capture complex interactions among different brain regions. This structure enhances the AT-DGNN’s capacity for integrating temporal features effectively.
@@ II-A Multi-head Attention Mechanism
Each self-attention head comprises three fundamental components: queries $Q$, keys $K$, and values $V$. These elements facilitate the computation of attention scores that influence the weighting of the values. The process begins with normalizing the input $X_{w}$ via a layer normalization (LayerNorm):
$\begin{cases}q_{ht}=W^{Q}\textit{LayerNorm}(X_{w,t})\\ k_{ht}=W^{K}\textit{LayerNorm}(X_{w,t})\\ v_{ht}=W^{V}\textit{LayerNorm}(X_{w,t})\end{cases}$ (1)
Here, $q_{ht}$, $k_{ht}$, and $v_{ht}$ denote the queries, keys, and values at time $t$ for head $h$, derived from the normalized input. The matrices $W^{Q}$, $W^{K}$, and $W^{V}$ belong to $\mathbb{R}^{d\times d_{H}}$, where $d_{H}$ represents the dimension of each attention head. This normalization facilitates the efficient computation of attention scores.
The attention context vector $c_{ht}$ for each head is computed as a weighted sum of the values, with weights $\alpha_{htt^{\prime}}$ determined by the scaled dot-product attention mechanism. The alignment scores $e_{htt^{\prime}}$ are calculated as:
@@
Here, $t^{\prime}$ denotes the different time steps considered. The softmax function converts these scores into a probability distribution, reflecting the importance of each value vector:
$c_{ht}=\sum_{t^{\prime}}\alpha_{htt^{\prime}}v_{ht^{\prime}}$ (3)
where $\alpha_{htt^{\prime}}$ represents the normalized alignment scores, indicating the importance of each value vector $v_{ht^{\prime}}$ in forming the context vector $c_{ht}$ for time step $t$. This context vector $c_{ht}$ captures the most relevant information from the values, highlighting key features at each time step.
@@ II-B Graph Neural Networks
GNNs represent a significant advancement in the domain of neural networks, designed explicitly to process graph-structured data. Unlike CNNs, GNNs excel in capturing the intricate relationships and dependencies among nodes within a graph through processes of aggregation and propagation of information across local neighbors [17]. A typical graph is denoted as $G=(V,E)$, where $V$ symbolizes the set of nodes and $E$ the set of edges. Each node $v_{i}\in V$ and edge $e_{ij}=(v_{i},v_{j})\in E$ can be respectively associated with a node and an edge in the graph. The adjacency matrix $A$ is configured as an $n\times n$ matrix where $A_{ij}=1$ if $e_{ij}\in E$ and $A_{ij}=0$ otherwise. Node attributes are represented by $X$, where $X\in\mathbb{R}^{n\times d}$, with each $x_{i}\in\mathbb{R}^{d}$ denoting the feature vector of node $v_{i}$.
@@
$\textit{h}_{t,i}^{(k+1)}=\sigma\left(\sum_{j\in\mathcal{N}(v_{i})}\alpha_{t,ij}^{(k)}W^{(k)}\textit{h}_{t,j}^{(k)}+\textit{b}^{(k)}\right)$ (4)
Here, $\mathcal{N}(v_{i})$ represents the neighbors of node $v_{i}$, $\alpha_{t,ij}^{(k)}$ are the attention coefficients indicating the significance of the features of neighbor $j$ to node $i$, and $W^{(k)}$ and $b^{(k)}$ are trainable parameters of the $k$-th layer, with $\sigma(\cdot)$ being a non-linear activation function.
@@ II-C Graph Neural Networks for EEG
In 2019, Song et al. [18] first introduced a novel DGCNN specifically for multichannel EEG emotion recognition, employing an adjacency matrix to dynamically model EEG channel relationships and enhance feature discrimination. This method proved to be superior to existing approaches. Subsequently, advancements in GNNs have been increasingly applied to emotion recognition. On the SEED dataset, Bao et al. [19] integrated a multi-layer GNN with a style-reconfigurable CNN, while Asadzadeh et al. [20] further improved the DGCNN by incorporating Bayesian signal recovery techniques, both achieving enhanced performance. For the DEAP dataset, Zhao et al. [21] developed coherence-based node clustering and multi-pooling techniques for GCNs. Additionally, Zhang et al. [22] enhanced GCN feature extraction by using Granger causality analysis and differential entropy to construct more informative connectivity matrices.
@@ III Methodology
In this section, we introduce the detailed architecture of AT-DGNN from the three aspects of data preprocessing, feature extraction and graph-based learning in turn, and the overall structure of AT-DGNN is shown in Fig. 1.
@@ III-A EEG Data Preprocessing
The initial sampling rate of the MEEG dataset is 1000 Hz. To align with standard EEG datasets typically sampled at 128 Hz or 200 Hz, and to mitigate issues such as high artifact noise and overlapping interference signals, the data are downsampled to 200 Hz. This preprocessing phase involves the use of a band-pass filter with a range of 1 to 50 Hz to enhance the signal quality by minimizing interference. Subsequently, the EEG signals are segmented into five distinct frequency bands for feature extraction: Delta (1-4 Hz), Theta (4-8 Hz), Alpha (8-14 Hz), Beta (14-31 Hz), and Gamma (31-50 Hz) [23].
@@ III-B1 Temporal Learner
A temporal learner layer employing multiscale 1D temporal kernels ($T$ kernels) is used to directly extract dynamic temporal representations from EEG data $X_{i}\in\mathbb{R}^{E\times T}$, where $E$ denotes the number of EEG electrodes, and $T$ is the sample length. These kernels obviate the need for manually extracted features. The $i$th-kernel’s length ($S_{T}^{i}$), dictated by the sampling frequency ($f_{s}$) of EEG data and scaling coefficients ($\alpha^{i}$), is given by [24]:
@@
EEG data processed through these layers yield dynamic time-frequency representations. An average pooling (AvgPool) layer, acting as a window function, calculates the averaged power across shorter segments. The logarithmic activation as described by [10] is applied to improve the performance. The output from each layer $i$, denoted as $Z^{i}_{\textit{temp}}$, is formulated as:
$Z^{i}_{\textit{temp}}=\Phi_{\textit{log}}(\textit{AvgPool}(\Phi_{\textit{sqr}}(\mathcal{F}_{\textit{Conv1-D}}(X_{i},S^{i}_{T}))))$ (6)
Here, $\Phi_{\textit{log}}(\cdot)$ is the logarithmic activation function, $\Phi_{\textit{sqr}}(\cdot)$ represents the square function, and $\mathcal{F}_{\textit{Conv1-D}}(\cdot)$ signifies the 1D convolution operation.
@@
$Z_{T}=f_{\textit{bn}}(Z^{1}_{\textit{temp}},Z^{2}_{\textit{temp}},Z^{3}_{\textit{temp}})$ (7)
where $f_{\textit{bn}}(\cdot)$ denotes the batch normalization operation.
@@ III-B2 Sliding Window Segmentation
Following the temporal learner, a convolution-based sliding window technique as [25] is applied to segment the EEG time series $X$ into multiple windows. This method integrates sliding window segmentation with convolutional operations, significantly reducing computational overhead by allowing convolutions to be executed once across all windows. This approach not only enhances data augmentation but also accelerates processing through parallel computation. The time series $X$ is segmented into windows $X_{w}\in\mathbb{R}^{B\times C\times W}$ using a window of length $W$ and stride $S$, where $w=1,\ldots,n$ represents the window index and $n$ is the total number of windows. Each window $X_{w}$ is then processed by subsequent attention and temporal convolution blocks. The number of windows $n$ is calculated as:
$n=\left\lfloor\frac{\textit{length}-\textit{window\_size}}{\textit{stride}}\right\rfloor+1$ (8)
Here, length, window_size, and stride denote the total length of the time series, the length of each window, and the stride between windows, respectively.